OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0).

jigglypuff314 (jigglypuff314):

first find the slope: \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] when given points (x1 , y1) and (x2, y2) so plug in x1 = -6 , y1 = 4 , x2 = 2 , y2 = 0 and simplify Then point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1) you would plug in what you got for slope then plug in a given point...